A raft is 2.35 m wide by 1.25 m long by 0.365 m deep. What is the bouyant force on the raft if only 0.125 of the raft remains above water?

1
Expert's answer
2021-11-04T10:17:32-0400

"F=mg=\rho g V=\rho g ab\cdot(1-0.125)c="


"=1000\cdot9.8\cdot2.35\cdot1.25\cdot(1-0.125)\cdot0.365=9194\ (N)" . Answer